OpenStudy (anonymous):

how do you find the slope and y-intercept of 12x-6y=30

OpenStudy (anonymous):

break it down into slope intercept form \[y = mx +b\]

OpenStudy (anonymous):

Move it to the form y=mx+b. 12x-6y=30 -12x -12x -6y=-12x+30 /(-6) /(-6) y=2x-5

OpenStudy (anonymous):

Thank you!